OpenStudy (anonymous):

A company’s new product is sold for $16.50 per unit. Product sales in the United States earn $400,000 and product sales in Europe earn $230,000. How many units does it sell altogether in the United States and Europe? a. 13,939 units c. 24,242 units b. 38,181 units d. 381,181 units

OpenStudy (anonymous):

To solve for United States:\[\frac{ 400,000 }{ 16.50 } = a\] To solve for Europe: \[\frac{ 230,000 }{ 16.50 } = b\] To solve for the total:\[a + b = total\] Understand?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Make sure to close the question unless you have more.

OpenStudy (anonymous):

is the answer b

OpenStudy (anonymous):

@Nibby

OpenStudy (anonymous):

I'm not sure why you have the answers in that order, but yes, the answer is B, 38,181.

OpenStudy (anonymous):

sorry is was supposed to be like this a c b d

OpenStudy (anonymous):

It's fine, but yes, your answer is B.

OpenStudy (anonymous):

thx